4. A researcher is interested in whether participating in sports positively influences self-esteem in young girls. She identifies a group of girls who have not played sports before but are now planning to begin participating in organized sports. The researcher gives them a 50-item self-esteem inventory before they begin playing sports and administers the same test again after 6 months of playing sports. The self-esteem inventory is measured on an interval scale, with higher numbers indicating higher self-esteem. In addition, scores on the inventory are normally distributed. The scores follow.

Before=X

After=Y

D = |X-Y|

D - D ̅

(D-D ̅ )^2

44

46

2

0.5

0.25

40

41

1

-0.5

0.25

39

41

2

0.5

0.25

46

47

1

-0.5

0.25

42

43

1

-0.5

0.25

43

45

2

0.5

0.25

Sum=9

Sum=1.5

21pOBRaTbjNSAVboO7Yns6IBKyt07ffgdsLdgfUK = Mean of the difference scores = 9/6 = 1.5

SD = 6xmT0b76BXmBzXSEEpOwAAAAAElFTkSuQmCC = 0.548

7g1i43ZcruuPTQAAAABJRU5ErkJggg===HqJvoLbHeyMfP2wyPmgf9gs9+3PD5zTgOvfo9q5o

A) Should H0 be rejected?

B) If significant, compute and interpret the effect size.

C) If significant, draw a graph representing the data.

D) Determine the 95% confidence interval.

I already did the initial analysis, but I am stuck on the parts listed above.
Answer #1

a)

µD = Mean of X – Mean of Y

alpha = 0.05

Null and Alternate Hypothesis

H0: µD = 0 (No impact)

Ha: µD < 0 (Positive impact of playing sports)

Test Statistic

t = Mean/ Std Dev = -1.5/0.548 = -2.74

P-value = TDIST(2.74,5,1) = 0.020

Result

Since the p-value is less than 0.05, we reject the null hypothesis in favour of alternate hypothesis ie participating in sports positively influences self-esteem in young girls.

b)

SDpooled = {(s12+s22)/n1+n2-2}1/2 = 1.15

Effect Size = (xQ - xM )/ SDpooled = 1.5/1.15 = 1.30

The high value of effect size tells us that there is meaningful difference between X and Y.

c)

Comparison between X andY 49 47 45 43 41 39 37 35 Before-XAfter-Y

d)

95% CI of difference of means Mean +/- 1.96 * Std Dev = -1.5+/- 1.96*0.548 = {-2.57, -0.43}